QUIZ #11: CALCULUS 1A (Stankova) Wednesday, April 14, 2004 Section 10:00–11:00 (Voight)

Name:

Please complete the following problem(s) in the space provided. You may not use a calculator. You will have 15 minutes to complete the quiz.

Please include all relevant intermediate calculations and explain your work when appropriate.

Problem 1. Find the antiderivative F of f that satisfies the given condition.

f(x) = 4 − sec^{2}(2x), F (π) = 0.

QUIZ #11: CALCULUS 1A (Stankova) Wednesday, April 14, 2004 Section 11:00–12:00 (Voight)

Name:

Please complete the following problem(s) in the space provided. You may not use a calculator. You will have 15 minutes to complete the quiz.

Please include all relevant intermediate calculations and explain your work when appropriate.

Problem 1. Find f, if

f′′(x) = sin x + x^{−1/2}.